# TransferTOD: A Generalizable Chinese Multi-Domain Task-Oriented Dialogue System with Transfer Capabilities

### v1

@@ Abstract
Task-oriented dialogue (TOD) systems aim to efficiently handle task-oriented conversations, including information gathering. How to utilize ToD accurately, efficiently and effectively for information gathering has always been a critical and challenging task. Recent studies have demonstrated that Large Language Models (LLMs) excel in dialogue, instruction generation, and reasoning, and can significantly enhance the performance of TOD through fine-tuning. However, current datasets primarily cater to user-led systems and are limited to predefined specific scenarios and slots, thereby necessitating improvements in the proactiveness, diversity, and capabilities of TOD. In this study, we present a detailed multi-domain task-oriented data construction process for conversations, and a Chinese dialogue dataset generated based on this process, TransferTOD, which authentically simulates human-machine dialogues in 30 popular life service scenarios. Leveraging this dataset, we trained a TransferTOD-7B model using full-parameter fine-tuning, showcasing notable abilities in slot filling and questioning. Our work has demonstrated its strong generalization capabilities in various downstream scenarios, significantly enhancing both data utilization efficiency and system performance. The data is released in https://github.com/KongLongGeFDU/TransferTOD.
@@
Ming Zhang1,   Caishuang Huang1,   Yilong Wu1 ,   Shichun Liu1,   Huiyuan Zheng1, Yurui Dong1,   Yujiong Shen1,   Shihan Dou1,   Jun Zhao1,   Junjie Ye1, Qi Zhang1,   Tao Gui2,   Xuanjing Huang1 1 School of Computer Science, Fudan University 2 Institute of Modern Languages and Linguistics, Fudan University mingzhang23@m.fudan.edu.cn qz.fudan.edu.cn
@@ 1 Introduction
The Task-Oriented Dialogue System (TOD) is a human-computer interaction system aims to aid users in accomplishing specific tasks or acquiring particular information, which has found extensive use in daily life and commercial applications. At present, TOD systems have displayed the capability to adapt effectively to diverse tasks, domains, and user behaviors. Nonetheless, they continue to encounter various challenges related to generality, deep understanding, proactive questioning, and other aspects.
@@
Figure 1: The system will ask the user for one or more slot values that are ‘none’, and then identify and update the corresponding field(s) based on the user’s response until all slots are filled. For instance, if the system inquires about the user’s age, and the user replies with ‘23’, the model will update the slot form ‘none’ to 23.
To gather the necessary information, the system must proactively ask questions or guide users to provide the required information for filling specific slots, known as slot filling (SF) Rosset et al. (2011). Although various approaches Devlin et al. (2019); Liu et al. (2019); Henderson et al. (2020) has been explored to maxmize data efficiency (e.g. transfer learning and fine-tuning), traditional SF methods still rely on expert-labeled data Fuisz et al. (2022), which is costly and inefficient, and are limited to predetermined scenarios and time periods. These methods cannot be generalized to more general scenarios easily (certain methods even rely on external databases Zhou et al. (2018); Tian et al. (2022); Zou et al. (2021)), and it is difficult to ensure accurate, real, and diverse responses to user needs. Furthermore, existing datasets primarily revolve around user-driven systems, where the focus is on constructing systems that primarily respond to user inquiries and requests Wen et al. (2017); Budzianowski et al. (2020); Zhu et al. (2020). Recently, Large Language Models (LLMs) have exhibited promising performance in dialogue participation, instruction generation, and zero-shot reasoning Zhang et al. (2023), which brought new ideas to solving the above problems. Research has confirmed that fine-tuned LLMs on dialogue corpora of different sizes can achieve enhanced performance across diverse tasks, domains, and even languages Du et al. (2021); Touvron et al. (2023). Hence, we can use LLMs to drive TOD and solve some problems that were difficult to solve in the small model era.
In the paper, we introduce TransferTOD: a multi-domain, task-oriented Chinese dialogue dataset encompassing more complex and diverse dialogue tasks, and simulating more realistic conversation scenarios. Inspired by real-world questionnaire-style information-gathering scenarios, TransferTOD facilitates interactions between users and systems to assist in information acquisition and record updating. The dataset includes 35965 turns of statements and 5460 dialogues across 30 popular life service scenarios, providing researchers with a more challenging and practically significant dataset. Considering potential human errors and the variability of the Chinese context in practical applications, we have incorporated perturbed data and data polished through various methods into the dataset.
By selecting appropriate base models and fine-tuning methods, we have successfully demonstrated that training the TransferTOD-7B model using our dataset can achieve high accuracy. This model not only can proactively ask users for missing slots and accurately fill them based on their answers but also performs efficiently in guiding fluency and generating responses. Additionally, we evaluated the quality of the models in terms of slot filling ability and semantic accuracy in guiding user responses. The results indicate that our dataset can significantly improve model performance by handling noise, increasing question diversity, and optimizing language fluency.
@@
1. We construct a new dataset called TransferTOD for task-oriented dialogue generation in various lifestyle service scenarios. It consists of 30 scenarios with 5460 dialogues, and ablation experiments have demonstrated that this dataset exhibits good noise resistance, diversity, and fluency.
2. We present a comprehensive dataset construction pipeline with high generalizability and transferability, enabling fellow researchers to effectively apply the methodology for creating datasets across various languages or in multilingual contexts.
3. We have utilized TransferTOD as our SFT dataset and trained the TransferTOD-7B model through full-parameter fine-tuning, achieving better slot filling and questioning capabilities compared to GPT-4. Additionally, with appropriate secondary fine-tuning techniques, our model demonstrates superior performance in out-of-domain testing compared to GPT-3.5-Turbo fine-tuned with an equivalent amount of data.
@@ 2 Related Work
The performance of intelligent dialogue systems is profoundly influenced by the quality of the dialogue datasets, making dataset construction an active research area. Initial generations of task-oriented dialogue datasets often focused on a single task or even a single scenario, ATIS Hemphill et al. (1990), DSTC2 Henderson et al. (2014), WOZ2.0 Wen et al. (2017), etc. included. The emergence of these databases not only enhanced the conversational fluency of conversational agents but also made task completion through natural dialogues between machines and humans possible. Considering that user dialogues often involve domain transitions, datasets Multi-WoZ Budzianowski et al. (2020), CrossWoZ Zhu et al. (2020) etc. encompassing more scenes and larger volumes of data were subsequently proposed. However, these dialogues are user-led discussions on relevant topics, requiring a user to pose questions or set tasks for the dialogue agent to respond accordingly.
@@
Enhancing the performance and data utilization of TOD systems and strengthening their ability to understand specific tasks expressed by users remain hot research topics. To complete tasks and improve accuracy, Li et al. (2018) proposed an end-to-end neural dialogue system based on reinforcement learning. TOD gradually started to realize across tasks Peng et al. (2017), domains Hakkani-Tür et al. (2016), and even languages Wang et al. (2021). TOD-BERT Wu et al. (2020), MinTL Lin et al. (2020), Soloist Peng et al. (2021), etc. has been successively proposed improving the success rate of tasks. However, as task complexity increases, these methods still rely heavily on large-scale datasets and lack competitiveness in handling noise robustness.
@@ 3.1 DataSet
TransferTOD aims to construct a cross-disciplinary task-oriented information collection multi-turn dialogue dataset, encompassing tasks such as goal-oriented questioning, dialogue state maintenance, information collection, and parsing. Existing task-oriented Wizard of Oz (WoZ) datasets are typically user-driven systems with relatively single domains. Departing from scenarios in the real world where questionnaire-style information collection may occur, we have curated dialogues spanning 30 different domains. We have enhanced the data in terms of robustness, diversity, and fluency, ensuring that the data closely mirrors real-world situations.
@@ 4.1.3 Evaluation Metrics
For the out-of-domain test, we assess the model’s capabilities in two main aspects: slot filling ability and semantic accuracy during the phase of guiding user responses. To evaluate the slot filling ability, we employ F1 and Joint Accuracy, which are widely used in the TOD systems for slot extraction tasks. To evaluate the semantic accuracy of model-generated questions, we use a manual evaluation approach. Please refer to the appendix A.3 for details. It is worth noting that we use the Dialog Act F1 as the evaluation metric in this context. While the Dialog Act F1 and SlotF1 metrics appear to be calculated in the same way, they differ slightly in essence. For a detailed explanation of these subtle differences, please refer to the Appendix A.3.
@@ 4.2.1 Results on In-Domain Test
Table 2: Results of the in-domain test: The dialogue act F1 Score of each model, showing the accuracy of predicting the right dialogue acts from user utterance.
@@ 4.2.2 Results on Out-Of-Domain Test
Table 3: Result of out-of-domain: The Joint Accuracy and Slot F1 Score of each model, showing the accuracy of predicting the right dialogue state and slot-value pairs respectively.
@@ 4.3.1 Secondary Fine-Tuning
In this section, we primarily discuss our experiments on performing secondary fine-tuning on TransferTOD-7B. The objective was to simulate enhancing our model’s slot filling and question-asking capabilities in external scenarios using a small subset of downstream scenario data. We fine-tuned GPT-3.5-Turbo [^5] https://platform.openai.com/docs/guides/fine-tuning as our baseline and conducted fine-tuning with 50, 100, and 200 pieces of data across three out-of-domain scenarios, respectively. The remaining data served as the test set for this experiment.
@@ 4.3.2 Result
Table 4: Result of Secondary Fine-Tune: The Joint Accuracy and Slot F1 Score of each model, showing the accuracy of predicting the right dialogue state and slot-value pairs respectively. "OTD" stands for Original Train Data which is used in fine-tuning the TransferTOD-7B. "200$\times$4" in Num.ScenarioData represents that we took 200 ScenarioData and repeated it four times.
@@ 4.4 Ablation Studies
Based on the TransferTOD-v1, v2, v3, and v4 mentioned in 3.1, we trained models TransferTOD-7B-v1 to v4 individually. To ascertain the efficacy and trustworthiness of our data construction methodologies, we rigorously assessed their performance in terms of robustness, diversity and fluency. The method we employed, which combines GPT-based assessment with expert review, is a widely adopted approach for evaluating the language fluency of models Chang et al. (2024); Zhang et al. (2024a). For details on the GPT assessment instructions and the expert review process, please refer to the appendix C.2, E.2.
@@
Table 5: Result of Noise Injection: The Joint Accuracy and Slot F1 Score of each model, showing the accuracy of predicting the right dialogue state and slot-value pairs respectively.
@@
Figure 3: Results of comparative evaluation of TransferTOD-7B-v2 and TransferTOD-7B-v3.
@@
Table 6: Comparison of GPT and Expert Reviews for TransferTOD-7B-v3 and TransferTOD-7B-v4’s inquiring fluency. The table shows the proportion of high and low scoring questions in GPT and expert ratings, as well as the corresponding total score.
@@ 5 Conclusion
In summary, we have proposed a highly versatile data construction process that enhances the quality of task-oriented dialogue data for information gathering tasks. The models fine-tuned with this data exhibit strong generalization capabilities, performing well in out-of-domain scenarios.
@@ A.2 Implementation Details
Table 7: Hyper-Parameters adopted when training TransferTOD-7B
@@ Appendix B Templates for script-generated data
Table 9: The prompt for rewriting the question
@@ Appendix C Prompts
Table 10: The prompt for rewriting the question
@@
Table 11: The prompt for rewriting the answer
@@ C.2 Prompt GPT-4 to Evaluate the Results
The prompt in Table 12 is used when using GPT-4 to conduct comparative evaluation of diversity in ablation experiments, while the prompt in Table 13 is used when scoring the fluency of model questions in ablation experiments with GPT-4.
@@
Table 12: The prompt used when using GPT-4 to conduct comparative evaluation of diversity in ablation experiments.
@@
Table 13: The prompt used when scoring the fluency of model questions in ablation experiments using GPT-4.
@@ E.1 Experts in Constructing Datasets
During the dataset construction phase, we relied on 5 students from our institute to participate in this work as human experts. These students possessed basic computer knowledge and coding skills, which enabled them to perform the task effectively.
@@ E.2 Experts in Ablation Experiment
Table 14: Result of out-of-domain with the setting of In-Context Learning: The Joint Accuracy and Slot F1 Score of each model, showing the accuracy of predicting the right dialogue state and slot-value pairs respectively.

### v2

@@ Abstract
Task-oriented dialogue (TOD) systems aim to efficiently handle task-oriented conversations, including information collection. How to utilize TOD accurately, efficiently and effectively for information collection has always been a critical and challenging task. Recent studies have demonstrated that Large Language Models (LLMs) excel in dialogue, instruction generation, and reasoning, and can significantly enhance the performance of TOD through fine-tuning. However, current datasets primarily cater to user-led systems and are limited to predefined specific scenarios and slots, thereby necessitating improvements in the proactiveness, diversity, and capabilities of TOD. In this study, we present a detailed multi-domain task-oriented data construction process for conversations, and a Chinese dialogue dataset generated based on this process, TransferTOD, which authentically simulates human-computer dialogues in 30 popular life service scenarios. Leveraging this dataset, we trained a model using full-parameter fine-tuning called TransferTOD-7B, showcasing notable abilities in slot filling and questioning. Our work has demonstrated its strong generalization capabilities in various downstream scenarios, significantly enhancing both data utilization efficiency and system performance. The data is released in https://github.com/KongLongGeFDU/TransferTOD.
@@
Ming Zhang1thanks:   Equal Contribution.,   Caishuang Huang1∗,   Yilong Wu1∗ ,   Shichun Liu1∗,   Huiyuan Zheng1∗, Yurui Dong1,   Yujiong Shen1,   Shihan Dou1,   Jun Zhao1,   Junjie Ye1, Qi Zhang1thanks:   Corresponding Author.,   Tao Gui2,   Xuanjing Huang1 1 School of Computer Science, Fudan University 2 Institute of Modern Languages and Linguistics, Fudan University mingzhang23@m.fudan.edu.cn qz@fudan.edu.cn
@@ 1 Introduction
The Task-Oriented Dialogue (TOD) system is a human-computer interaction system aims to aid users in accomplishing specific tasks or acquiring particular information, which has found extensive use in daily life and commercial applications. At present, TOD systems have displayed the capability to adapt effectively to diverse tasks, domains, and user behaviors. To collect the necessary information, the system must proactively ask questions or guide users to provide the required information for filling specific slots, known as slot filling (SF) Rosset et al. (2011).
@@
Figure 1: In information collection scenaries, the system will ask the user for one or more slot values that are ‘none’, and then identify and update the corresponding field(s) based on the user’s response until all slots are filled. For instance, if the system inquires about the user’s age, and the user replies with ‘23’, the system will update the slot form ‘none’ to 23.
Although various approaches has been explored to enhance data efficiency (e.g. transfer learning and fine-tuning) Devlin et al. (2019); Liu et al. (2019); Henderson et al. (2020) , traditional SF methods still rely on expert-labeled data Fuisz et al. (2022), which is costly and inefficient, and are limited to predetermined scenarios and time periods. Moreover, existing slot filling datasets mostly serve user-driven dialogue systems, typically constrained to specific scenarios and slots predetermined for constructing systems that respond to user inquiries and requests Wen et al. (2017); Budzianowski et al. (2020); Zhu et al. (2020). In system-driven information collection scenarios, the slots and values that need to be filled are often unfamiliar with the training set, leading to a significant decline in the accuracy of slot filling. Additionally, the system struggles to ask precise questions based on the filled slots to guide user responses, thereby completing the next slot-filling action. Traditional TOD systems face significant challenges in this task.
Recently, Large Language Models (LLMs) have exhibited promising performance in dialogue participation, instruction generation, and zero-shot reasoning Zhang et al. (2023), which brought new ideas to solving the above problems. Research has confirmed that fine-tuned LLMs on dialogue corpora of different sizes can achieve enhanced performance across diverse tasks, domains, and even languages Du et al. (2021); Touvron et al. (2023). Hence, we can use LLMs to drive TOD and solve some problems that were difficult to solve in the small model era.
In this paper, inspired by real-world customer information collection scenarios as shown in the Figure 1, we proposed a dataset called TransferTOD. This multi-domain, task-oriented Chinese dialogue dataset is designed to handle complex and diverse tasks, simulating realistic conversation scenarios. To account for potential human errors and practical applications in the Chinese context, we implemented a four-step dataset construction pipeline, including noise injection and language polishing. This process resulted in a robust dataset containing 35965 turns and 5460 dialogues across 30 life service scenarios, offering a valuable resource for research. Using this dataset, we trained the TransferTOD-7B model with appropriate base models and fine-tuning methods. The model can proactively ask users for missing information, accurately fill slots, and efficiently guide fluent response generation. Our evaluations of slot-filling ability and semantic accuracy demonstrated that the dataset significantly improves model’s performance by handling noise, increasing question diversity, and optimizing language fluency.
@@
1. We propose a comprehensive four-step pipeline with high generalizability, allowing researchers to create new TOD datasets across different languages or in multilingual contexts through similar methods.
2. We construct a new dataset called TransferTOD for task-oriented dialogue generation in various lifestyle service scenarios, containing 5460 dialogues across 30 scenarios. Ablation experiments have demonstrated that this dataset exhibits good noise resistance, diversity, and fluency.
3. We utilize TransferTOD to fine-tune the TransferTOD-7B model, achieving superior slot-filling and questioning capabilities compared to GPT-4. With appropriate secondary fine-tuning, our model demonstrates superior performance in out-of-domain test compared to GPT-3.5-Turbo fine-tuned with an equivalent amount of data.
@@ 2 Related Work
The performance of intelligent dialogue systems is profoundly influenced by the quality of the dialogue datasets, making dataset construction an active research area. Initial generations of task-oriented dialogue datasets often focused on a single task or even a single scenario, ATIS Hemphill et al. (1990), DSTC2 Henderson et al. (2014), WOZ2.0 Wen et al. (2017), etc. included.
The emergence of these databases not only enhanced the conversational fluency of conversational agents but also made task completion through natural dialogues between machines and humans possible. Considering that user dialogues often involve domain transitions, datasets Multi-WoZ Budzianowski et al. (2020), CrossWoZ Zhu et al. (2020) etc. encompassing more scenes and larger volumes of data were subsequently proposed.
However, these dialogues are user-led discussions on relevant topics, requiring a user to pose questions or set tasks for the dialogue agent to respond accordingly.
@@
Enhancing the performance and data utilization of TOD systems and strengthening their ability to understand specific tasks expressed by users remain hot research topics. To complete tasks and improve accuracy, Li et al. Li et al. (2018) proposed an end-to-end neural dialogue system based on reinforcement learning. TOD gradually started to realize across tasks Peng et al. (2017), domains Hakkani-Tür et al. (2016), and even languages Wang et al. (2021). TOD-BERT Wu et al. (2020), MinTL Lin et al. (2020), Soloist Peng et al. (2021), etc. has been successively proposed improving the success rate of tasks. However, as task complexity increases, these methods still rely heavily on large-scale datasets and lack competitiveness in handling noise robustness.
@@ 3.1 Dataset
TransferTOD aims to construct a cross-disciplinary task-oriented customer information collection multi-turn dialogue dataset, encompassing tasks such as goal-oriented questioning, dialogue state maintenance, information collection, and parsing. Existing task-oriented Wizard of Oz (WoZ) datasets are typically user-driven systems with relatively single domains. Departing from scenarios in the real world where customer information collection dialogue may occur, we have curated dialogues spanning 30 different domains. We have enhanced the data in terms of robustness, diversity, and fluency, ensuring that the data closely mirrors real-world situations.
@@ 4.1.3 Evaluation Metrics
For the out-of-domain test, we assess the model’s capabilities in two main aspects: slot filling ability and semantic accuracy during the phase of guiding user responses. To evaluate the slot filling ability, we employ F1 and Joint Accuracy, which are widely used in the TOD systems for slot extraction tasks. To evaluate the semantic accuracy of model-generated questions, we use a manual evaluation approach. Please refer to the appendix A.3 for details.
It is worth noting that we use the Dialog Act F1 as the evaluation metric in this context. While the Dialog Act F1 and SlotF1 metrics appear to be calculated in the same way, they differ slightly in essence. For a detailed explanation of these subtle differences, please also refer to the appendix A.3.
@@ 4.2.1 Results on In-Domain Test
Table 2: Results on in-domain test: The Dialogue Act F1 Score of each model, showing the accuracy of predicting the right dialogue acts from user utterance.
@@ 4.2.2 Results on Out-of-Domain Test
Table 3: Results on out-of-domain test: The Joint Accuracy and Slot F1 Score of each model, showing the accuracy of predicting the right dialogue state and slot-value pairs respectively.
@@ 4.3.1 Secondary Fine-Tuning
In this section, we primarily discuss our experiments on performing secondary fine-tuning on TransferTOD-7B. The objective was to simulate enhancing our model’s slot filling and question-asking capabilities in external scenarios using a small subset of downstream scenario data. We fine-tuned GPT-3.5-Turbo [^5] https://platform.openai.com/docs/guides/fine-tuning as our baseline and conducted fine-tuning with 50, 100 and 200 pieces of data across three out-of-domain scenarios, respectively. The remaining data served as the test set for this experiment.
@@ 4.3.2 Result
Table 4: Result of Secondary Fine-Tuning on out-of-domain test: The Joint Accuracy and Slot F1 Score of each model, showing the accuracy of predicting the right dialogue state and slot-value pairs respectively. "OTD" stands for Original Train Data which is used in fine-tuning the TransferTOD-7B. "200$\times$4" in Num.ScenarioData represents that we took 200 ScenarioData and repeated it four times.
@@ 4.4 Ablation Studies
Based on the TransferTOD-v1, v2, v3, and v4 mentioned in 3.1, we trained models TransferTOD-7B-v1 to v4 individually. To ascertain the efficacy and trustworthiness of our data construction methodologies, we rigorously assessed their performance in terms of robustness, diversity and fluency. The method we employed, which combines GPT-based assessment with expert review, is a widely adopted approach for evaluating the language fluency of models Chang et al. (2024); Zhang et al. (2024a). For details on the GPT assessment instructions and the expert review process, please refer to the appendix C.2 and E.2.
@@
Table 5: Results on Noise Injection: The Joint Accuracy and Slot F1 Score of each model, showing the accuracy of predicting the right dialogue state and slot-value pairs respectively.
@@
Figure 3: Results on Language Augmentation: Different color blocks represent the winning rate of two version models in terms of diversity comparison.
@@
Table 6: The results on Fluency Enhancement: The table shows the proportion of high and low scoring questions in GPT and expert ratings, as well as the corresponding total Ask_Flu score.
@@ 5 Conclusion
In summary, we have proposed a highly versatile data construction process that enhances the quality of task-oriented dialogue data for information collection tasks. The models fine-tuned with this data exhibit strong generalization capabilities, performing well in out-of-domain scenarios.
@@ A.2 Implementation Details
Table 7: Hyper-Parameters adopted when training TransferTOD-7B.
@@ Appendix B Templates for script-generated data
Table 9: An example of a question and answer template.
@@ Appendix C Prompts
Table 10: The prompt for rewriting the question.
@@
Table 11: The prompt for rewriting the answer.
@@ C.2 Prompt GPT-4 to Evaluate the Results
The prompt in Table 12 is used when using GPT-4 to conduct comparative evaluation of diversity in ablation experiments, while the prompt in Table 13 is used when using GPT-4 to score the fluency (Ask_Flu) of model questions in ablation experiments.
@@
Table 12: The prompt for using GPT-4 to conduct comparative evaluation of diversity in ablation experiments.
@@
Table 13: The prompt for using GPT-4 to score the fluency (Ask_Flu) of model’s questions in ablation experiments.
@@ E.1 Experts in Constructing Datasets
During the dataset construction phase, we relied on 5 students from our institute to participate in this work as human experts. These students possessed good computer knowledge and coding skills, which enabled them to perform the task effectively.
@@ E.2 Experts in Ablation Experiment
Table 14: Result of out-of-domain test with the setting of In-Context Learning: The Joint Accuracy and Slot F1 Score of each model, showing the accuracy of predicting the right dialogue state and slot-value pairs respectively.